Find an equation of the line having the given slope and containing the given point.
m= 5/6, (1, -7)
The equation of the line is y=____
(Simplify your answer. Use integers or fractions for any numbers in the expression)

Answer by Alan3354(69443)

Find an equation of the line having the given slope and containing the given point.
m= 5/6, (1, -7)
The equation of the line is y=____
----------------
Use the slope-intercept equn
y-y1 = m*(x-x1)
y-(-7) = (5/6)*(x-1)
y+7 = (5/6)x -5/6
y = (5x/6) - 47/6